Grade 12th passAlgebra

Find the square of the following numbers.
35..........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

(35)2
= (30+5 )2
= (30)2 + (5)2 + 2×30×5 [Since, (a+b)2 = a2+b2 +2ab]
=900+25+300
= 1225

Thanks